### v1

# Benchmarking Algorithms for Federated Domain Generalization

Abstract
While prior domain generalization (DG) benchmarks consider train-test dataset heterogeneity, we evaluate Federated DG which introduces federated learning (FL) specific challenges. Additionally, we explore domain-based heterogeneity in clients’ local datasets—a realistic Federated DG scenario. Prior Federated DG evaluations are limited in terms of the number or heterogeneity of clients and dataset diversity. To address this gap, we propose an Federated DG benchmark methodology that enables control of the number and heterogeneity of clients and provides metrics for dataset difficulty. We then apply our methodology to evaluate 13 Federated DG methods, which include centralized DG methods adapted to the FL context, FL methods that handle client heterogeneity, and methods designed specifically for Federated DG. Our results suggest that despite some progress, there remain significant performance gaps in Federated DG particularly when evaluating with a large number of clients, high client heterogeneity, or more realistic datasets. Please check our extendable benchmark code here: https://github.com/inouye-lab/FedDG_Benchmark.
1 Introduction
Domain generalization (DG) (Blanchard et al., 2011) formalizes a special case of train-test heterogeneity in which the training algorithm has access to data from multiple source domains but the ultimate goal is to perform well on data from an unseen test domain—i.e., a type of out-of-distribution generalization instead of the standard in-distribution generalization. While most prior DG work focuses on centralized algorithms, another natural context is federated learning (FL) (Konečnỳ et al., 2016), which is a distributed machine learning context that assumes each client or device owns a local dataset. These local datasets could exhibit heterogeneity, which we call client heterogeneity (e.g., class imbalance between clients). Although train-test heterogeneity (in DG) and client heterogeneity (in FL) are independent concepts, both could be naturally defined in terms of domain datasets. For example, suppose a network of hospitals aimed to use FL to train a model to predict a disease from medical images. Because the equipment is different across hospitals, it is natural to assume that each hospital contains data from different domains or environments (or possibly a mixture of domains if it is a large hospital)—this is a case of domain-based client heterogeneity. Yet, the trained model should be robust to changes in equipment within a hospital or to deployment in a new hospital that joins the network—both are cases of domain-based train-test heterogeneity. The interaction between these types of heterogeneity produces new algorithmic and theoretic challenges yet it may also produce new insights and capabilities. Solutions to Federated DG with both types of heterogeneity could increase the robustness and usefulness of FL approaches because the assumptions more naturally align with real-world scenarios rather than assuming the datasets are i.i.d. This could enable training on partial datasets, increase robustness of models to benign spatial and temporal shifts, and reduce the need for retraining.
In the centralized regime, various approaches have been proposed for DG, including feature selection, feature augmentation, etc. Most of these methods are not applicable in the FL regime which poses unique challenges. In the FL regime, client heterogeneity has long been considered a statistical challenge since FedAvg (McMahan et al., 2017), where it experimentally shows that FedAVG effectively mitigate some client heterogeneity. There are many other extensions based on the FedAvg framework tackling the heterogeneity among clients in FL, for example using variance reduction method (Karimireddy et al., 2020). An alternative setup in FL, known as the personalized setting, aims to learn personalized models for different clients to tackle heterogeneity, for example Hanzely et al. (2020). However, none of these works consider model robustness under domain shift between training and testing data. Recently, a few works in the FL regime tackling DG (Liu et al., 2021; Zhang et al., 2021; Nguyen et al., 2022; Tenison et al., 2022) have been proposed, however their evaluations are limited in the following senses: 1) The evaluation datasets are limited in the number and diversity of domains. 2) The evaluations are restricted to the case when the number of clients is equal to the number of domains, which may be an unrealistic assumption (e.g., a hospital that has multiple imaging centers or a device that is used in multiple locations). The case when clients number might be massive are of both theoretical and application interests. 3) None of the works consider the influence of the effect of the number of communication rounds. We provide an overview of the tasks inTable 1, considering both the heterogeneity between training and testing datasets (standard vs. domain generalization) and among clients (domain client heterogeneity). While some studies have addressed the standard supervised learning task, there is a need for a fair evaluation to understand the behavior of domain generalization algorithms in the FL context under those new challenges.
There are several benchmark datasets available for evaluating domain generalization (DG) methods in the centralized setting. These benchmarks, such as DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021), provide multiple datasets that are suitable for assessing the performance of DG algorithms. However, they did not explicitly consider the unique challenges that arise in the federated learning (FL) setting. On the other hand, there are also benchmarks specifically designed for FL. For instance, the LEAF benchmark (Caldas et al., 2018) provides a standardized framework for evaluating FL algorithms. It includes several datasets from various domains and allows researchers to assess the performance of their algorithms in a realistic FL scenario. Another benchmark for FL is PFLBench (Chen et al., 2022), which focuses on evaluating personalized FL methods. PFLBench provides 12 datasets containing various applications. Though these FL-based benchmarks consider statistical heterogeneity, they fail to consider the DG task adequately. Moreover, the level of statistical heterogeneity present in these datasets is insufficient for proper DG evaluation. In summary, DG benchmarks do not consider FL challenges, and FL benchmarks do not consider DG challenges.
Major contributions: We develop a benchmark methodology for evaluating Federated DG with various client heterogeneity contexts and diverse datasets, and we evaluate representative Federated DG approaches with this methodology. 1) We propose the first Federated DG benchmark methodology including four important dimensions of the experimental setting (see Section 3). 2) We propose a standardized definition of domain-based client heterogeneity that is unique to the FL context and interpolates between domain homogeneity and domain separation (see Fig. 1) while controlling class imbalance. In particular, we develop a novel method to split any dataset with domain labels across any number of clients (see subsection 3.2). 3) We compare three broad approaches to Federated DG: centralized DG methods naïvely adapted to FL setting, FL methods developed for client heterogeneity (e.g., class imbalance), and recent methods specifically designed for Federated DG. Our results indicate that there still exist significant gaps and open research directions in Federated DG. 4) We will release an extendable open-source library for evaluating Federated DG methods.
Notation
[…]
2 Approaches to Federated Domain Generalization
We first briefly review the DG problem, extend to the Federated DG problem, and explain domain-based client heterogeneity. Then, we discuss three categories of Federated DG approaches.
2.1 Problem Background and Setup
Domain generalization (train-test heterogeneity)
Unlike standard ML which assumes the train and test are independent and identically distributed (i.i.d.), the ultimate goal of DG is to minimize the average or worst-case loss of the test domain distributions when only samples from the train domain distributions are given. Formally, given a set of train domain distributions $\{p_{d}:d\in\mathcal{D}_{\text{train}}\}$, minimize the average or worst case loss over test domain distributions, i.e.,
$\min_{\theta}{\textstyle\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{d\in\mathcal{D}_{\text{test}}}}\mathcal{L}(\theta;p_{d})\quad\text{or}\quad\min_{\theta}\max_{d\in\mathcal{D}_{\text{test}}}\mathcal{L}(\theta;p_{d})\,,$ (1)
where $\mathcal{L}(\theta;p_{d})=\mathbb{E}_{(x,y)\sim p_{d}}[\ell(x,y;\theta)]$ where $\ell$ is a per-sample loss function such as squared or cross-entropy loss. The key challenge in DG is that the train and test domain distributions are disjoint, i.e., $\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{test}}=\emptyset$, and thus, the method must be able to generalize beyond the train domain distributions to perform well on the test domain distributions. The naïve approach is to simply ignore the domains and perform empirical risk minimization (ERM) on all the training data—which is actually a challenging baseline to outperform in practice (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021).
Federated DG
Federated DG adds a layer of complexity because now the domain distribution samples are not centrally located. Instead, each FL client can update their local model based only on their local client distribution $p_{c}$ and then pass their local parameters $\theta_{c}$ to a central server, which will aggregate the local models and broadcast the model to all clients. The FL problem can be abstracted as follows:
$\forall c\in[C],\,\,\underbrace{\theta_{c}=\texttt{Optimize}(\mathcal{L}(\theta;p_{c}),\theta_{\text{init}}=\theta_{\text{global}})}_{\text{Locally optimize given local distribution $p_{c}$}}\quad\text{and}\quad\underbrace{\theta_{\text{global}}=\texttt{Aggregate}(\theta_{1},\theta_{2},\cdots,\theta_{C})}_{\text{Aggregate client model parameters on server}}\,,$
where the client distributions may be homogeneous (i.e., $\forall(c,c^{\prime}),p_{c}=p_{c^{\prime}}$) or heterogeneous (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$), Optimize minimizes an objective $\mathcal{L}(\theta;p_{c})$ initialized at $\theta_{\text{init}}$, and Aggregate aggregates the client model parameters, where the most common aggregator is simply a (weighted) average of the client parameters, which corresponds to FedAvg (McMahan et al., 2017).
Domain-based client heterogeneity
While client heterogeneity (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$) is often expressed as label imbalance, i.e., $p_{c}(y)\neq p_{c^{\prime}}(y)$, we make a domain-based client heterogeneity assumption that each client distribution is a (different) mixture of train domain distributions, i.e., $p_{c}(x,y)=\sum_{d\in\mathcal{D}_{\text{train}}}w_{c,d}\,\,p_{d}(x,y)$ where $w_{c,d}$ is the weight of the $d$-th domain for the $c$-th client. At one extreme, FL with i.i.d. data would be equivalent to the mixture proportions being the same across all clients, i.e., $\forall c,c^{\prime},d,w_{c,d}=w_{c^{\prime},d}$, which we call the homogeneous setting. On the other extreme, if the number of clients and number of domains are equal ($D=C$), the domain weights could be disjoint between clients such that each client “owns” one or more domains, i.e., for any domain $d\in[D],$ $w_{c,d}>0\Leftrightarrow w_{c^{\prime},d}=0,\forall c^{\prime}\neq c$, which we call domain separation (in subsection 3.2, we extend the domain separation to the case $D\neq C$). Finally, we use $\lambda$ to denote an interpolation parameter between these two extremes: homogeneous case $\lambda=1$ and domain separation case $\lambda=0$. We give details of an explicit algorithm to use in practice for splitting datasets for evaluation in subsection 3.2 which covers two extremes and interpolation cases $\lambda\in(0,1)$. Table 1 summarizes the train-test heterogeneity in DG and the domain-based client heterogeneity from the FL context, where we focus on Federated DG.
Table 1: Domain-based Federated DG considers the domain heterogeneity both among the clients’ local datasets (rows) and between the train and test datasets (columns). This paper focuses on the DG setting (right column).
Between train and test datasets (train-test heterogeneity)
Standard supervised learning Domain generalization (our focus)
Among clients Homogeneous ($\lambda=1$) FL with homogeneous clients Federated DG with homogeneous clients
Heterogeneous ($10<\lambda<$ 1) FL with heterogeneous clients Federated DG with heterogeneous clients
Domain separation ($\lambda=0$) FL with domain separation Federated DG with domain separation
Overview of Federated DG methods
In this benchmark study, we explore three categories of Federated DG methods: DG methods originally designed for the centralized setting, FL methods specifically tailored to handle client heterogeneity, and methods specifically designed for Federated DG. To provide a comprehensive evaluation, we assess the performance of several representative methods from each of these categories and compare to vanilla FedAvg (McMahan et al., 2017) with ERM loss, where any heterogeneity is simply ignored. These methods are selected based on their prominence and potential effectiveness for Federated DG. To ensure a diverse range of evaluation scenarios, we conduct experiments on various datasets and under various heterogeneity settings. Our results provide insights into their relative performance and suitability for different Federated DG contexts. This benchmark study aims to offer a broad overview of the current popular methods, enabling researchers and practitioners to make informed decisions when tackling Federated DG.
2.2 Centralized DG methods adapted to the FL setting
The first natural choice is directly migrating the DG methods from the centralized regime. To adapt those methods, we simply run the centralized DG method at each client locally with their own local dataset (see subsection 3.2 for how the local datasets are created), and then compute an average of model parameters at each communication round (see next paragraph). This approach is straightforward for the homogeneous ($\lambda=1$) and heterogeneous ($\lambda=0.1$) settings where each client has data from all training domains—albeit quite imbalanced for $\lambda=0.1$. This can be seen as biased updates at each client based on biased local data. In the domain separation case $\lambda=0$, (i.e., if all clients only have one primary domain, i.e., $\forall k,|P_{k}|=1$), this simple approach cannot be applied because centralized DG methods require data from at least two domains. In fact, these centralized DG methods degenerate to vanilla FedAvg if there is only one domain per client. Extending these methods to the case where all clients only have one domain without violating the FL constraints is an interesting direction for future work.
A predominant and effective centralized DG approach is through representation learning, including domain-invariant representation learning by learning domain-invariant features via either kernel methods (Muandet et al., 2013; Ghifary et al., 2016) or invariant risk minimization (Arjovsky et al., 2019; Krueger et al., 2020), and domain adversarial neural networks Sun and Saenko (2016), and Li et al. (2018b) which explicitly tries to align the feature representation distribution. Besides invariant representation, there are other methods the general learning strategy to promote the generalization ability gradient operation Shi et al. (2021); Rame et al. (2022), which tries to learn generalized representations by directly operating on gradients. Other approaches include distributionally robust optimization Sagawa et al. (2019), which learns the worst-case distribution scenario of training domains; and meta-learning Finn et al. (2017); Li et al. (2018a), which is based on the learning-to-learn mechanism to learn general knowledge by constructing meta-learning tasks to simulate domain shift. We selected IRM (Arjovsky et al., 2019), Fish (Shi et al., 2021), MMD (Li et al., 2018b), Mixup (Zhang et al., 2017), DeepCoral (Sun and Saenko, 2016), and GroupDRO (Sagawa et al., 2019) from this category for their representative mechanisms.
2.3 FL methods tackling client heterogeneity
Another line of research in FL aims to guarantee convergence even under client heterogeneity, but these FL-based methods still assume the train and test datasets do not shift (i.e., they do not explicitly tackle train-test heterogeneity of the domain generalization task). The empirical observation of the statistical challenge in federated learning when local data is non-IID was first made by Zhao et al. (2018). Several subsequent works have analyzed client heterogeneity by assuming bounded gradients (Yu et al., 2019; Basu et al., 2019; Wang et al., 2019; Li et al., 2019) or bounded gradient dissimilarity (Li et al., 2020), and additionally assuming bounded Hessian dissimilarity (Karimireddy et al., 2020; Khaled et al., 2020; Liang et al., 2019). From this category, we selected FedProx (Li et al., 2020), which addresses statistical heterogeneity by adding a proximal term to the local subproblem, constraining local updates to be closer to the initial global model. Scaffold (Karimireddy et al., 2020) utilizes variance reduction to account for client heterogeneity.
2.4 FL methods designed for Federated DG
Limited research has focused explicitly on solving the Federated DG by design. FedDG Liu et al. (2021) introduced a specific FL paradigm for medical image classification, which involves sharing the amplitude spectrum of images among local clients, violating the privacy protocol. Another approach, FedADG (Zhang et al., 2021), utilizes a generative adversarial network (GAN) framework in FL, where each client contains four models: a featurizer, a classifier, a generator, and a discriminator. FedADG first trains the featurizer and classifier using empirical loss and then trains the generator and discriminator using the GAN approach. However, this method requires training four models simultaneously and tuning numerous hyperparameters, making convergence challenging. A novel aggregation method called Federated Gradient Masking Averaging (FedGMA) (Tenison et al., 2022) aims to enhance generalization across clients and the global model. FedGMA prioritizes gradient components aligned with the dominant direction across clients while assigning less importance to inconsistent components. FedSR (Nguyen et al., 2022) proposes a simple algorithm that uses two locally-computable regularizers for domain generalization. Given the limited literature on solving domain generalization (DG) in the federated learning (FL) setting, we selected all the aforementioned algorithms.
3 Benchmark Methodology
In this study, we aimed to conduct a comprehensive evaluation of the Federated DG task by considering four distinct dimensions of the problem setup. We evaluated a total of $13$ methods, encompassing three different types of approaches.
3.1 Four evaluation dimensions
(1) Domain-based client heterogeneity and (2) number of clients
Previous studies on Federated DG have often focused on domain separation client heterogeneity where the number of clients equals the number of training domains, i.e., $C=D$. However, this excludes evaluation of the homogeneous and partially heterogeneous settings and restricts the number of clients to the number of training domains $D$. In particular, many pseudo-realistic datasets, such as those in DomainBed (Gulrajani and Lopez-Paz, 2020), which consist training domains $D\leq 5$, which limits the evaluation of methods under scenarios with a large number of clients. Conversely, when using realistic domain datasets that have many domains, such as those in WILDS (Koh et al., 2021), most current methods perform poorly under this extremely challenging setting. By introducing the domain split method discussed in subsection 3.2, we can explore various levels of client heterogeneity and relax the assumption that $C=D$ so that we can leverage both pseudo-realistic and realistic datasets and evaluate methods at an appropriate difficulty level.
(3) Dataset difficulty and (4) dataset type
While most Federated DG work focuses on standard image-based datasets, we evaluate methods across a broad range of dataset difficulty (ranging from easy pseudo-realistic datasets to very challenging realistic datasets) and two types of datasets ($3$ image datasets and $2$ text datasets). This ensures that we can fully understand the performance of each method across a wide range of scenarios. Further details can be found in Section subsection 3.3.
3.2 Domain-based client heterogeneity by splitting DG datasets
For evaluation dimensions (1) and (2) above, we need to control over the amount of client heterogeneity, ranging from homogeneous to domain separation, and we need to control over the number of clients, which may be smaller or larger than than the number of domains (i.e., $D\neq C$). We propose a way to split any DG dataset into any number of clients that allows explicit control of the amount of heterogeneity through the $\lambda$ hyperparameter which attempts to balance the number of samples per client. An illustration of our domain splitting procedure can be found in the following Fig. 1.
Figure 1: Distinct color refers to distinct domain data, and $\lambda$ is the domain balancing parameter. (a): Train-test domain heterogeneity. (b): Domain split when $C\leq D$ and domain split when $C>D.$ (c): domain split when domain number exceeds client number, i.e., $C\leq D;$ homogeneous ($\lambda=1$), heterogeneous ($\lambda=0.1$), and separation ($\lambda=0$). See Algorithm 1 for domain split when $C>D.$
Algorithm 1 DomainSplit Algorithm
Input Number of clients $C$, heterogeneity parameter $\lambda\in[0,1]$, and samples from each domain $(\mathcal{S}_{1},\cdots,\mathcal{S}_{D})$, where $n_{1},\dots,n_{D}$ denote the number of samples per domain and w.l.o.g. are assumed to be in descending order, i.e., $n_{1}\geq n_{2}\geq\dots\geq n_{D}$.
  // Divide domain indices across clients
  if $C\leq D$ then
     $\forall c,\mathcal{D}_{c}\leftarrow\emptyset$
     for $d=1,2,\dots,D$ do
        // Find client with fewest samples
        $c^{*}\in{\textstyle\operatorname*{arg\,min\,}\limits_{c\in[C]}\,\sum_{d^{\prime}\in\mathcal{D}_{c}}}n_{d^{\prime}}$
        $\mathcal{D}_{c^{*}}\leftarrow\mathcal{D}_{c^{*}}\cup\{d\}$
     end for
  else if $C>D$ then
     $\forall c\in\{1,2,\dots,D\},\,\mathcal{D}_{c}\leftarrow\{c\}$
     for $c=D+1,\dots,C$ do
        // Find on average largest domain to split
        $d^{*}\in\operatorname*{arg\,max\,}\limits_{d\in[D]}\frac{n_{d}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]}$
        $\mathcal{D}_{c}\leftarrow\{d^{*}\}$
     end for
  end if
  // Split samples across the clients
  $\forall c,\mathcal{S}_{c}\leftarrow\emptyset$
  for $c\in[C],d\in[D]$ do
     $n_{d,c}(\lambda)=\lambda\frac{n_{d}}{C}+(1-\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]\cdot n_{d}}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c}]}$
     $\mathcal{S}^{\prime}\leftarrow\texttt{SampleWOReplacement}(\mathcal{S}_{d},n_{d,c})$
     $\mathcal{S}_{d}\leftarrow\mathcal{S}_{d}\setminus S^{\prime}$// Remove from domain
     $\mathcal{S}_{c}\leftarrow\mathcal{S}_{c}\cup S^{\prime}$// Append to client dataset
  end for
Output $(\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{C})$
In Algorithm 1, we provide a concrete algorithm for splitting samples from multiple domains across an arbitrary number of clients $C$ while controlling the amount of domain-based client heterogeneity via $\lambda$, ranging from homogeneous clients to domain separation. Our algorithm has two main steps. In Step 1, we assign “primary” domain indices $\mathcal{D}_{c}\subseteq[D]$ to each client $c\in[C]$ depending on $C$ and $D$. If $C\leq D$, the domains are sorted in descending order according to number of sample size and are iteratively assigned to the client $c^{*}$ which currently has the smallest number of training samples (denoted by $\sum_{d^{\prime}\in\mathcal{D}_{c^{*}}}n_{d^{\prime}}$). In this case, the algorithm ensures that no client shares domains with the others but otherwise attempts to balance the total number of training samples between clients. If $C>D$, we first assign the domains one by one to the first $D$ clients. Then, starting from client $c=D+1$, we iteratively assign the current on average largest domain $d^{*}$ to $c,$ accounting for the fact that the domain may already shared by different clients—notationally, on average is represented through dividing by $\sum_{c^{\prime}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]$. In this case, some clients may share one domain, but no client holds two domains simultaneously while otherwise attempting to balance the number of samples across clients as much as possible.
In Step 2, we define the sample counts for each client and domain, denoted $n_{c,d}(\lambda)$ based on the balancing parameter $\lambda\in[0,1]:$ $n_{d,c}(\lambda)=\lambda\frac{n_{d}}{C}+(1\!\!-\!\!\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c}]}n_{d}\,,$ where rounding to integers is carefully handled when not perfectly divisible and where $\mathbbm{1}[\cdot]$ is the indicator function. This is simply a convex combination between homogeneous clients ($\lambda=1$) and domain separation ($\lambda=0$). Given the number of samples per client per domain, we simply sample without replacement from the corresponding domain datasets and build up the client datasets.
3.3 Dataset Selection and Dataset Difficulty Metrics
To ensure a comprehensive evaluation of current methods across different difficulty levels, we have curated a range of datasets with varying complexities. We define two dataset metrics to measure the dataset difficulty with respect to the DG task and with respect to the FL context. For DG difficulty, we compute $R_{\text{DG}},$ the ratio of the ERM performance with and without samples from the test domain (i.e., the former is able to “cheat” by seeing part of test domain samples during training). For FL difficulty, we attempt to isolate the FL effect by computing $R_{\text{FL}}(\lambda),$ the ratio of ERM-based FedAvg $\lambda$ client heterogeneity over centralized ERM on in-domain test samples. These dataset difficulty metrics can be formalized as follows:
$R_{\text{DG}}$ $\triangleq\frac{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}^{\prime}_{\text{DG-test}})}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}}\cup{\mathcal{S}^{\prime\prime}_{\text{DG-test}}},\mathcal{S}^{\prime}_{\text{DG-test}})}$
$R_{\text{FL}}(\lambda)$ $\triangleq\frac{\texttt{FedAvg\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}};\lambda)}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}})},$
where ERM_Perf is the performance of ERM using the first argument as training and the second for test, FedAvg_Perf is similar but with the client heterogeneity parameter $\lambda$, $\mathcal{S}_{\text{DG-train}}$ denotes samples from the train domains $\mathcal{D}_{\text{train}},$ $\mathcal{S}_{\text{DG-test}}$ denotes samples from the test domains $\mathcal{D}_{\text{test}},$ and $\mathcal{S}^{\prime}_{\text{DG-test}}$ and $\mathcal{S}^{\prime\prime}_{\text{DG-test}}$ are $20\%,80\%$ split respectively of $\mathcal{S}_{\text{DG-test}}$. For $R_{\text{FL}}(\lambda)$, we use $\mathcal{S}_{\text{IN-test}}$ (test samples from the train domains) instead of $\mathcal{S}_{\text{DG-test}}$ to isolate the FL effect from DG effect. We choose five datasets in our benchmark: FEMNIST from Caldas et al. (2018), PACS from Gulrajani and Lopez-Paz (2020), and IWildCam, CivilComments and Py150 from Koh et al. (2021). We summarize the statistics and difficulty metrics in Table 2 and provide the rationale of selecting these datasets in the Appendix.
Table 2: Summary of selected datasets as well as their difficulty metric, where $n$ is the number of samples, $D$ is the number of domains, $C$ is the number of clients used.
Dataset Statistics Difficulty Metric
$n$ $D$ $C$ $R_{\text{DG}}$ $R_{\text{FL}}(\lambda)$
$\lambda=1$ $\lambda=0.1$ $\lambda=0$
FEMNIST $737036$ $2586$ $100$ $1.000$ $0.980$ $0.981$ $0.981$
PACS $9991$ $2$ $100$ $0.960$ $1.000$ $1.000$ $1.000$
IWildCam $203029$ $243$ $243$ $0.449$ $0.682$ $0.525$ $0.525$
CivilComments $448000$ $16$ $100$ $0.984$ $0.618$ $0.533$ $0.532$
Py150 $150000$ $8421$ $100$ $0.969$ $0.999$ $0.998$ $0.998$
4 Benchmark Experimental Results
In this section, we report the performance of $13$ representative methods from three lines of research on $5$ different datasets, where the FEMINIST results are provided in the appendix given the DG task simplicity (i.e., $R_{\text{DG}}\approx 1$). For each dataset, we fix the total computation and communication rounds for different methods for a fair comparison. We then select the model based on the held-out-domain validation set. After training, we choose the model according to the early-stopping at the communication round which achieves the best held-out-domain performance, and finally we evaluate the performance on the test-domain in Table 3 and Table 4. See Appendix for detailed hyperparameters choices. We also include the results according to in-domain validation early-stopping in the Appendix but the trends are similar. We make the following remarks on the main results from Table 3 and Table 4.
Remark 4.1.
FedAvg with an ERM objective is a strong baseline. Simple FedAvg with an ERM objective is a strong baseline that is challenging to beat across datasets (except for CivilComments), similar to the centralized case stated in DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021). We recommend always including FedAvg as a baseline in all future evaluations.
Table 3: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
PACS ($D=2$) IWildCam ($D=243$)
$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.8379$ $0.8766$ $0.9144$ $\mathit{0.9377}$ $0.2727$ $0.1762$ $0.1241$ $0.0707$
DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $\mathit{0.3134}$ $0.0718$ $0.0924$
DeepCoral $\bf 0.9216$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
GroupDRO $0.9060$ $\bf 0.9395$ $\bf 0.9437$ $0.2531$ $0.0530$ $0.0908$
FL FedProx - $0.8826$ $0.8407$ $0.8892$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
Scaffold - $0.8215$ $0.8796$ $0.9239$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
FDG FedDG - $\mathit{0.9234}$ $\mathit{0.9275}$ $\mathbf{0.9521}$ $0.1445$ $0.1191$ $0.0503$
FedADG - $0.1892$ $0.0592$ $0.0598$ $0.0049$ $0.0049$ $0.0049$
FedSR - $0.1246$ $0.1263$ $0.1257$ $0.0056$ $0.0056$ $0.0056$
FedGMA - $0.8820$ $0.8467$ $0.8446$ $0.0106$ $0.0957$ $0.0106$
Table 4: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
CivilComments ($D=16$) Py150 ($D=5477$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $\mathit{0.6405}$
DG Adapted IRM $0.6408$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
Fish $\mathbf{0.6713}$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
MMD $\mathit{0.6520}$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $0.6565$
DeepCoral $0.5847$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
GroupDRO $0.6383$ $0.6028$ $0.4954$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
FDG FedSR - 0.3603 0.3389 0.3188 - 0.0915 0.0860 0.0587
FedGMA - 0.3341 0.3157 0.3195 - 0.6628 0.3737 0.3576
Remark 4.2.
Most centralized DG methods degrade in the FL setting. For image datasets, the DG methods adapted to the FL setting (except for GroupDRO) show significant degradation in performance compared to their centralized counterparts as can be seen when comparing the $C=1$ column to the $C>1$ columns in Table 3. Further, degradation can be seen in PACS when moving from the homogeneous client setting ($\lambda=1$) to the heterogeneous client setting ($\lambda=0.1$).
Remark 4.3.
FL methods tackling client heterogeneity perform surprisingly well on the Federated DG task. FedProx and Scaffold, which were designed for client heterogeneity but not specifically for the DG task, perform quite well in the Federated DG setting and even perform the best for IWildCam and Py150. Thus, we suggest including these methods in future Federated DG evaluations and suggest that future works could focus on combining the strengths of these methods and DG methods.
Remark 4.4.
The performance of real-world data significantly degrades as $\lambda$ decreases. This can be seen from IWildCam and Py150. While it is challenging and expensive to run models for IWildCam and Py150, they show the largest differences between methods and demonstrates the real-world challenge of Federated DG. We suggest including IWildCam and Py150 in most future DG evaluations given their unique nature across datasets.
Remark 4.5.
A diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG. This conclusion is inspired from the following two observations. 1) When comparing FL methods, we notice opposite trends between PACS and all other datasets (IWildCam, CivilComments, Py150) when increasing client heterogeneity based on $\lambda$. For PACS, the FL methods (except for FedADG and FedSR) surprisingly seem to improve when increasing client heterogeneity (i.e., $\lambda=0$). However, for other datasets, the expected trend exists such that increasing client heterogeneity produces worse performance. From this, we recommend that using PACS alone might be a misleading or at least incomplete evaluation of Federated DG methods. 2) Centralized DG methods adapted to FL perform the best on CivilComments. This unexpected result suggests that DG methods in the centralized regime may be able to better accommodate subpopulation shift, which is the special kind of shift that is exhibited in CivilComments. Both of these observations emphasize that a diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG tasks.
Additional DG challenges from FL.
For further understanding, we explore some additional questions on the PACS dataset because it is the most common and computationally feasible to train many different models. Specifically, we explore how the number of clients, amount of communication (i.e., the number of server aggregations) in the federated setup, and client heterogeneity affects the performance of various methods. The figures and detailed analysis are provided in the Appendix but we highlight two remarks here.
Remark 4.6.
The number of clients $C$ strongly affects on the overall performance. The DG performance drops from $90\%$ to $50\%$ or even 10% when varying $C$ from $1$ to $200$. We strongly recommend future Federated DG evaluations to consider larger number of clients such as $50$ to $100$ rather than only a very small number of clients. It indicates that there exist significant unresolved challenges in the domain of federated domain generalization when dealing with a large number of clients, which have not been adequately explored. In particular, FedADG and FedSR seem to be sensitive to the number of clients because they perform poorly when $C\geq 10$, while in the original papers they were only evaluated on a few clients $C=3$ (and we reproduce the good results with such $C$ in the Appendix).
Remark 4.7.
The number of communications does not monotonically affect DG performance. We observe an interesting implicit regularization phenomena in the FL context: different methods achieve their best performance when the communication rounds is relatively low. In particular, for PACS, this communication round is $10$ (while fixing the total amount of computations). This is surprising as in contrast for in-domain task, FL suggests more communications leads to better performance (Stich, 2018, Theorem 5.1). Further theoretical investigation of the dependence of DG accuracy on the communications rounds, and its possible relation to implicit regularization via early stopping is an interesting area for future work.
5 Conclusion and Discussion
We first build a systematic benchmark methodology for evaluating Federated DG that includes novel methodologies for splitting the data across clients and evaluation of dataset difficulty. We then evaluate $13$ representative methods from three relevant lines of research. Our evaluation shows that Federated DG is still unsolved and significant gaps remain between centralized DG performance and Federated DG performance. Therefore, Federated DG is ripe for additional research. Based on our evaluation and observation, here are some recommendations and suggestions for future work in Federated DG.
Recommendations for future evaluations of Federated DG.
  1. 1.
    Stronger Baselines - FedAvg should always be included because it is a strong baseline (Remark 4.1). FL methods designed for client heterogeneity (though not necessarily DG) should also be included given their strong performance (Remark 4.3).
  2. 2.
    Realistic Datasets - Federated DG methods should be evaluated on more realistic datasets. In particular, FEMNIST and PACS behave quite differently with respect to client heterogeneity than the more realistic WILDS datasets (Remark 4.4). Thus, we recommend including both IWildCam and Py150 in future evaluations of Federated DG. CivilComments may be useful for evaluating realistic subpopulation shift (Remark 4.5).
  3. 3.
    Large Number of Heterogeneous Clients - Evaluations should include scenarios with a large number of clients with varying degrees of heterogeneity. A large number of clients poses unique challenges for most methods (Remark 4.6). Additionally, the heterogeneity of clients is more realistic and plays a significant role in evaluation (Remark 4.4).
Suggestions for future work in Federated DG.
  1. 1.
    Handle Domain Separation Case - The domain separation scenario ($\lambda=0$) limits the exchange of information between domains, i.e., a client may only have data from a single domain. Centralized approaches cannot be adapted to this setting and current methods still struggle under this realistic client heterogeneity setting.
  2. 2.
    Increase Convergence Rate - In most benchmark experiments (except for PACS, which is an easy dataset), we have observed slow convergence, particularly on challenging real-world datasets such as IWildCam and Py150. Moreover, we have noticed even slower convergence when client heterogeneity is high (i.e., when $\lambda$ is small). In addition, we have found that federated methods designed to improve convergence actually performed quite well even though they were not designed for the DG task (Remark 4.3). Therefore, it is crucial for future methods to improve the convergence rate of Federated DG approaches.
  3. 3.
    Investigate the Effect of Communication Frequency - Because increased communication frequency may actually hurt DG performance (Remark 4.7), further investigation into the regularization effect of early stopping and infrequent communications in FedAvg may yield important insights.
  4. 4.
    Understand the Effect of the Number of Clients - Given that the performance of some methods degrade very quickly when increasing the number of clients beyond a few (Remark 4.6), the community could benefit from an improved theoretic and empirical understanding of how the number of clients affects Federated DG in terms of convergence, computational requirements, sample complexity, and DG performance.
We hope this work provides a better foundation for future work in Federated DG and accelerates research progress.
Acknowledgements
This work was supported by Army Research Lab under Contract No. W911NF-2020-221. R.B. and D.I. also acknowledge support from NSF (IIS-2212097) and ONR (N00014-23-C-1016). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
Appendix
[…]
Reproducibility Statement
Code for reproduce the results is available at the following link:
https://github.com/inouye-lab/FedDG_Benchmark.
We include detailed documentation in using our code to reproduce the results throughout the paper. We also provide documentation in adding new algorithm’s DG evaluation in the FL context.
Appendix A Datasets and Difficulty Metric
A.1 Dataset Introduction
In this section, we introduce the datasets we used in our experiments, and the split method we used to build heterogeneous datasets in the training and testing phase as well as the heterogeneous local training datasets among clients in the FL.
FEMNIST
It is an FL prototyping image dataset of handwritten digits and characters each users created as a natural domains, widely used for evaluation for client heterogeneity in FL. Though it contain many training domains, it lacks significant distribution shifts across domains ($R_{\text{DG}}=1$), and considered as easy compared to other datasets.
PACS
It is an image dataset for domain generalization. It consists of four domains, namely Photo (1,670 images), Art Painting (2,048 images), Cartoon (2,344 images), and Sketch (3,929 images). This task requires learning the classification task on a set of objects by learning on totally different renditions. $R_{\text{DG}}=0.960$ makes it a easy dataset as domain generalization in our setting. Notice that choosing different domain as test domain might give us different $R_{\text{DG}}$.
IWildCam
It is a real-world image classification dataset based on wild animal camera traps around the world, where each camera represents a domain. It contains 243 training domains, 32 validation and 48 test domains. Usually people cares about rare speicies, thus we utilize the macro F1 score as the evaluation metric instead of standard accuracy, as recommended in the original dataset’s reference Koh et al. [2021]. The $R_{\text{DG}}=0.449$ makes it a very challenging dataset for domain generalization.
CivilComments
It is a real-world binary classification text-based dataset formed from the comments from different demographic groups of people, containing $8$ demographic group. The goal is to judge whether the text is malicious or not.
Py150
It is a real-world code-completion dataset which is challenging given massive domains $8421,$ where the domain is formed according to the repository author. The goal is to predict the next token given the context of previous tokens. We evaluate models by the accuracy on the class and method tokens.
A.2 Dataset Split Setup
For each dataset, we first split the dataset into $5$ categories, namely training dataset, in-domain validation dataset, in-domain test dataset, held-out validation dataset and test domain dataset. For FEMNIST and PACS dataset, we use Cartoon and Sketch as training domain, Art-painting as held-out domain, Painting as test domain. For training domain, we split $10\%,10\%$ of the total training domain datasets as in-domain validation dataset and in-domain test dataset respectively. For IWildCam, CivilComments and Py150, we directly apply the Wilds official splits.
Appendix B Benchmark Experimental Setting
B.1 Model Structure
In this benchmark, for image-based datasets, we use ResNet-50 He et al. [2016] dataset. For CivilComments and Py150 dataset, we choose DistilBERT Sanh et al. [2020] and CodeGPT Radford et al. [2019] respectively as recommended by Wilds.
B.2 Model Selection
We conduct held-out domain model selection with $4$ runs for each methods. The oracle model selection evaluates the model based on the performance on the held-out validation domain. The results are reported based on the best run.
B.3 Early Stopping
We conduct early stopping using the held-out validation dataset in our evaluation. For each dataset and method, We first run certain communication rounds, then we select the model parameters which achieves the best performance on the validation set. We report the held-out validation dataset in the main paper, and we report the results using the in-domain validation set in Appendix D.
B.4 Hyperparameters
In this section, we present the hyperparameters selected for the evaluation. We opted to communicate per epoch for all experiments. For PACS, we run totally $50$ communication rounds. For IWildCam, we run $50$ communication rounds. For CivilComments, we run $10$ communication rounds. For Py150, we run $10$ communication rounds, and for FEMNIST, we run $40$ communication rounds. Please refer to the Table 5 to review other hyperparameters.
Table 5: Hyperparameters used in the evaluation. c represents centralized learning regime, and f represents the federated learning regime
PACS IWildCam CivilComments Py150 FEMNIST
c f c f c f c f c f
Fed Avg Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
IRM Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty weight $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$
Penalty anneal iters $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$
Fish Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
meta learning rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Mixup Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\alpha$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$
MMD Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty Weight $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
Deep Coral Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty Weight $1$ $1$ $10$ $10$ $1$ $1$ $1$ $1$ $1$ $1$
Group DRO Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\eta$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$
Fed Prox Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\mu$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
Scaf fold Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
FedDG Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\lambda$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$
FedADG Classifier optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGDs
Classifier learning rate $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$
Gen optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
Gen learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
Disc optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
Disc learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
$\alpha$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$
FedSR Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\ell_{2}$ regularizer $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.001$ $0.001$ $0.01$ $0.01$
cmi regularizer $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
FedGMA Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Server step size $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
Mask threshold $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$
Appendix C Additional FL-specific challenges for domain generalization
As mentioned in Section 4, we also include some deeper exploration over the effect of number of clients and communication frequency, which are unique to the FL regime.
i) Massive number of clients: In this experiment, we explore the performance of different algorithms when the number of clients $K$ increases on PACS. We fix the communication rounds $50$ and the local number of epoch is $1$ (synchronizing the models every epoch). Fig. 2 plots the held-out DG test accuracy versus number of clients for different levels of data heterogeneity. The following comments are in order: given communication budget, 1) current domain generalization methods all degrade a lot after $C\geq 10$, while the performance ERM and FedDG maintain relatively unchanged as the clients number increases given communication budget. FedADG and FedSR are are sensitive to the clients number, and they both fail after $C\geq 20.$ 2) Even in the simplest homogeneous setting $\lambda=1,$ where each local client has i.i.d training data, current domain generalization methods IRM, FISH, Mixup, MMD, Coral, GroupDRO work poorly in the existence of large clients number, this means new methods are needed for DG in FL context when data are stored among massive number of clients.
Figure 2: PACS: Held-out DG test accuracy versus number of clients.
ii) Communication constraint: To show the effect of communication rounds on convergence, we plot the test accuracy versus communication rounds in Appendix Fig. 3. We fix the number of clients $C=100$ on PACS and decreases rounds of communication (together with increasing local epochs). That is, if the regime restricts the communication budget, then we increase its local computation $E$ to have the same the total computations. Therefore, the influence of communication on the performance is fair between algorithms because the total data pass is fixed. We observe that the total number of communications does not monotonically affect the DG performance. With decreasing number of total communication, most methods’ performance first increase then decrease. This might be an interesting implicit regularization phenomena in the FL context. Without discussing DG task, usually frequent communications lead to faster convergence. The relationship between DG performance and communications requires further exploration.
Figure 3: PACS: Held-out DG test accuracy vs. varying communications (resp. varying echoes ).
Appendix D Supplementary Results
In the main paper, we provide experiments results using held-out validation early stopping. Here we report the results using in-domain validation set. We also report the convergence curve for each methods on each dataset for reference. We observe that under most of cases, the held-out validation gives us a better model. Thus, we recommend using held-out validation set to perform early stopping.
More results on PACS and IWildCam dataset.
Here we both report the convergence curve using the DG accuracy of each method on PACS Fig. 4 and IWildCam Fig. 5 and the DG accuracy on PACS and IWildCam Table 6 using the in-domain validation. From the PACS convergence curve Fig. 4, we observe that with lower $\lambda$, the model is harder to converge. FedAvg ang FedDG converge faster than all other methods while FedADG and FedSR does not converge. From IWildCam convergence curve Fig. 5, we observe that all methods are struggling to converge. This is due to the main challenge first come from the $R_{\text{FL}}$ here. We thus observe that Federated methods dealing with heterogeneity clients achieves best performance. Especially when $\lambda$ is low. It is worth noticing that this does not mean DG in FL is solved, none of the methods even achieve the centralized ERM performance.
Figure 4: Convergence curve on PACS; total clients and training domains $(C,D)=(100,2);$ increasing domain heterogeneity from left to right: $\lambda=(1,0.1,0).$
Figure 5: Accuracy versus communication rounds for IWildCam; Total clients number $C=243$; increasing heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
Table 6: Test accuracy on PACS and IWildCam dataset with in-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
PACS ($D=2$) IWildCam ($D=243$)
$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.8389$ $0.8766$ $\mathit{0.9144}$ $\mathbf{0.8958}$ $\mathit{0.2727}$ $\mathbf{0.1762}$ $\mathbf{0.1241}$ $0.0707$
DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $0.3134$ $0.0718$ $0.0924$
DeepCoral $\mathbf{0.9216}$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
GroupDRO $0.9060$ $\mathbf{0.9395}$ $\mathbf{0.9437}$ $0.2531$ $0.0530$ $0.0908$
FL FedProx - $0.8325$ $0.8407$ $0.8568$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
Scaffold - $0.8215$ $0.8643$ $0.8787$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
FDG FedDG - $\mathit{0.8868}$ $0.8880$ $\mathit{0.8952}$ $0.1445$ $0.1023$ $0.0503$
FedADG - $0.0915$ $0.0732$ $0.0592$ $0.0049$ $0.0049$ $0.0049$
FedSR - $0.1180$ $0.1341$ $0.1138$ $0.0056$ $0.0056$ $0.0056$
FedGMA - $0.8820$ $0.8467$ $0.8180$ $0.0271$ $0.0259$ $0.0081$
More results on CivilComments and Py150 dataset.
Here we report the DG accuracy on CivilComments and Py150 Table 6 using the in-domain validation. We could observe that most of the methods perform well on CivilComments and Py150. This could be attributed to the use of pretrained models. Additionally, the utilization of pretrained models may explain the relatively high value of $R_{\text{DG}}$, as NLP pretrained models have already been exposed to various domains.
Table 7: Test accuracy on CivilComments and Py150 datasets with in-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
CivilComments ($D=16$) Py150 ($D=5477$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $0.6405$
DG Adapted IRM $\mathit{0.6213}$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
Fish $0.6038$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
MMD $0.5981$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $\mathit{0.6565}$
DeepCoral $0.5827$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
GroupDRO $\mathbf{0.6383}$ $0.6028$ $0.4804$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
FDG FedSR - $0.3603$ $0.3389$ $0.3188$ $0.0915$ $0.0860$ $0.0587$
FedGMA - $0.3341$ $0.3157$ $0.3195$ $0.6628$ $0.3737$ $0.3576$
Results on FEMNIST dataset.
As mentioned in the main paper, we include the FEMNIST dataset here as reference. We could observe from Table 8 that $\lambda$ does not influence the final DG accuracy, and either in-domain validation or held-out-domain validation does not affect the final DG accuracy. This indicates the lacking of the statistical heterogeneity across different domains. we observe that changing of $\lambda$ does not significantly affect the convergence. Most of which does not converge to the centralized counterpart’s performance, this is due to the challenge coming from large number of clients where $R_{\text{FL}}=0.980<1$.
Figure 6: Accuracy versus communication rounds for FEMNIST; total clients and training domains $(K,M)=(100,2586);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
Table 8: Test accuracy on FEMNIST dataset with in-domain validation and held-out validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
FEMNIST (In-Domain) FEMNIST (Held-Out-Domain)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$ $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$
DG Adapted IRM $0.8442$ $0.8303$ $0.8389$ $0.8328$ $0.8442$ $0.8303$ $0.8389$ $0.8328$
Fish $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$ $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$
Mixup $0.8337$ $0.8275$ $0.8168$ $0.8161$ $0.8337$ $0.8275$ $0.8168$ $0.8161$
MMD $0.8442$ $0.8430$ $0.8323$ $0.8293$ $0.8442$ $0.8430$ $0.8323$ $0.8293$
DeepCoral $0.8457$ $0.8403$ $0.8328$ $0.8513$ $0.8457$ $0.8403$ $0.8328$ $0.8513$
GroupDRO $0.8413$ $0.8447$ $0.8139$ $0.8052$ $0.8413$ $0.8447$ $0.8139$ $0.8052$
FL FedProx - $0.8448$ $0.8440$ $0.8425$ $0.8448$ $0.8440$ $0.8425$
Scaffold - $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$ $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$
FDG FedDG - $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$ $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$
FedADG - $0.0549$ $0.0557$ $0.0548$ $0.0549$ $0.0557$ $0.0548$
FedSR - $0.8320$ $0.8316$ $0.8306$ $0.8320$ $0.8316$ $0.8306$
FedGMA - $\mathit{0.8486}$ $0.8395$ $0.8371$ $0.8486$ $0.8395$ $0.8371$
Appendix E Gap Table
We list the gap table in Table 9 for summarizing the current DG algorithms performance gap w.r.t FedAvg-ERM in the FL context, in particular, positive means it outperforms FedAvg-ERM, negative means it is worse than FedAvg-ERM. It can be seen that in the on the simple dataset, the best DG migrated from centralized setting is better than FedAvg-ERM. In the domain separation case, no centralized DG algorithms can be adapted to it, and FDG methods performs comparably good in this setting. However, they fail in harder datasets. In the hardest setting, currently the Federated methods dealing with data heterogeneity performs the best. It is worth noting that while federated learning methods that address client heterogeneity perform better than other methods, they still fall short of achieving centralized empirical risk minimization (ERM). This highlights the need for future research and development of DG methods in the FL regime.
Table 9: Gap Table: Current progress in solving Federated DG with respect to FedAvg-ERM.
Domain Separation
Free of Data Leakage
PACS
IWildCam
CivilComments
Py150
Adapted centralized methods $+0.0629$ $-0.0981$ $+0.0421$ $-0.0062$
Federated methods $+0.0060$ $+0.0458$ $+0.0421$ $-0.0062$
FDG FedDG $+0.0468$ $-0.0775$
FedADG $-0.6874$ $-0.2171$
FedSR $-0.7520$ $-0.2164$ $-0.0326$ $-0.5912$
FedGMA $+0.0054$ $-0.1263$ $-0.0588$ $-0.0199$
Appendix F Training Time, Communication Rounds and local computation
In this section, we provide training time per communication in terms of the wall clock training time. Notice that for a fixed dataset, most of algorithms have similar training time comparing to FedAvg-ERM, where FedDG and FedADG are significantly more expensive.
Table 10: Wall-clock Training time per communication (unit: s).
PACS IWildCam CivilComments Py150 FEMNIST
$C=100$ $C=243$ $C=100$ $C=100$ $C=100$
B
FedAvg-ERM $143$ $6301$ $3958$ $6566$ $262$
DG Adapted IRM $147$ $6454$ $4085$ $7089$ $297$
Fish $148$ $7072$ $5483$ $7770$ $324$
Mixup $144$ $6294$ $264$
MMD $144$ $6663$ $4024$ $7603$ $287$
DeepCoral $144$ $6597$ $3901$ $7212$ $287$
GroupDRO $145$ $9311$ $4690$ $8121$ $307$
FL FedProx $169$ $6921$ $4502$ $6513$ $288$
Scaffold $167$ $6876$ $4421$ $6389$ $281$
FDG FedDG $352$ $32172$ $989$
FedADG $181$ $11094$ $907$
FedSR $151$ $7136$ $4403$ $8020$ $280$
FedGMA $143$ $6795$ $4525$ $6545$ $261$